# Federated Behavioural Planes: Explaining the Evolution of Client Behaviour in Federated Learning

### v1

@@ 5.3 Leveraging FBPs information
FBPs offer detailed insights into client behaviors, enabling the Federated Behavioural Shield to outperform existing state-of-the-art defense mechanisms. (Figure 4) Our comparative analysis demonstrates that Federated Behavioural Shields generally outperform traditional methods such as Krum, Median, and Trimmed-mean across various datasets including Breast, Diabetes, and small-MNIST. The only exception was in the scenario of the Inverted gradient on the small-MNIST dataset, where Trimmed-mean performed better. Overall, our method enhanced the performance up to 10 percentage points (pp) over Median—the most robust aggregation baseline—when the FL system is under Label-flipping attack and up to 16 pp when the system is not under attack on small-MNIST. The proposed approach surpasses even a FedAvg aggregation in absence of attackers on the Breast Cancer and on the small-MNIST datasets, under both normal and crafted-noise conditions. These results suggest that an aggregation strategy based on predictive performance or decision-making similarities is superior to methods that solely consider sample count. Unlike other baselines, the independence of our method from prior knowledge about attackers establishes it as a robust aggregation tool in both adversarial and non-adversarial settings. Contrary to previous studies [14, 28], we find that predictive performance alone does not reliably identify malicious clients in non-IID settings, as honest clients may also show consistently low performance, leading to potential misclassification in the aggregation process. Occasionally, this reliance on predictive performance can reduce our method’s accuracy compared to scenarios where only counterfactual information is used. For instance, under no-attack conditions, counterfactual information alone often offers more effective behavioural descriptors than combining it with predictive metrics. Further detailed analyses in Appendix A.7 indicate that using both descriptors generally yields better results than relying solely on counterfactual behaviour.
@@ 6.1 Related works
Federated Learning Client behaviours in FL have been analysed using integrated visual tools [31, 44, 45] or indirectly through methods such as robust aggregation [10, 11, 12, 13, 14, 28, 30, 32, 39, 46, 47, 48, 49, 50] and clustering-based aggregation [51, 52, 53, 54, 55, 56, 57]. These studies typically focus on similarities in model or gradient parameters [10, 11, 12, 13, 31, 39, 44, 45, 46, 47, 48, 49, 50, 53, 54, 55, 56, 57], under the premise that distinct client data distributions manifest as unique model parameters. Nevertheless, these methodologies might overlook valuable information about how client models behave with the data, particularly in terms of predictive performance. To mitigate this issue, previous studies adopted evaluation methods that utilise a clean validation set on the server to delineate clients’ behaviour. Common descriptors used in such cases are straightforward metrics such as accuracy, error rates, and loss [14, 31, 44, 45, 51, 52]. Despite their utility, these metrics might not be able to reveal behavioural patterns, such as those involved in the decision-making processes of models, which may suggest subtle similarities or manipulations. Notably, Wang et al. [44] introduced a post-hoc explainable approach, Grad-CAM [58], to explain client model behaviours during training. However, Grad-CAM is limited to CNN models and provides primarily qualitative visual insights, which cannot be easily automated and thus still require human intervention. Similarly, SentiNet [32] uses Grad-CAM to detect potentially malicious regions in input images, yet client behaviour assessments continue to rely primarily on evaluating predictive performance on manipulated and unmanipulated images. To the best of our knowledge, this work represents the first systematic attempt to formalise the evolving dynamics of clients in FL, showing how behavioural client trajectories affect the predictive performance and decision-making processes of the global model.
@@ 6.2 Limitations and future works
The primary constraint of our framework lies in its assumption that the server possesses a minimal validation set for querying client models. While this assumption is common across various methods, it can be mitigated as demonstrated in Wang et al. [44], by generating synthetic data points. To this end, our approach might already integrate a potential mechanism as differentiable counterfactual methods can be used to generate synthetic data [59]. A downside of counterfactual generators is that their concurrent training in FL systems introduces an overhead both in the optimisation objective and client-server communications. While in our experiments we did not observe a decreased predictive performance, this overhead may lead to higher computational demands and more advanced optimisation strategies in more complex scenarios. Future works may consider the extensive information available from FBPs to investigate other automatic strategies to optimise the learning process, such as the development of Clustered FL among clusters of clients and the fine-grained categorization of attack types. Incorporating additional behavioural planes may also enhance the specificity of the FBPs explanations.
@@ 6.3 Conclusions
In this work, we proposed Federated Behavioral Planes, a method to explain the dynamics of FL systems and client behaviours. This innovative method allows to visualise, track, and analyse client behaviours based on specific characteristics. Our focus was twofold: examining predictive performance by analysing prediction errors and investigating the decision-making process through counterfactual generation. The results of our experiments showed that Federated Behavioural Planes enable to track client behaviours over time, cluster similar clients, and identify clients’ contributions to the global model with respect to a specific descriptor. Based on Federated Behavioural Planes’ information, we introduced a novel robust aggregation mechanism that improves existing state-of-the-art methods by not requiring prior knowledge of the attacker. This work lays the foundation to explain the evolution of client behaviours, with the potential to enhance reliability and control over FL systems.
@@ A.1 Datasets
In our experiments, we employ four distinct datasets: a Synthetic dataset (tabular); the Breast Cancer Wisconsin dataset [36] (tabular); the Diabetes Health Indicator dataset [37] (tabular); and the MNIST dataset [38] (image). The Synthetic dataset consists of two features randomly extracted from a range of -5 to +5. As shown in Figure 5.a, we partition this feature space into $K$ slices, where $K$ equals the number of clients. Each slice corresponds to a specific client. We assign a label of 1 to all data points where $x_{1}>\alpha x_{2}$, effectively creating a linear decision boundary that varies with $\alpha$. For each client, we generally draw 1000 samples. This controlled setup allows us to precisely manipulate and visually understand the data distribution across different clients, as illustrated in Figure 5a. The Breast Cancer dataset contains data from 569 patients with 30 continuous variables derived from digitised images of a fine needle aspirate of a breast mass. The variables describe characteristics of the cell nuclei present in the image (e.g., radius, area, perimeter), aimed at predicting breast cancer. The Diabetes dataset comprises data from 70,692 patients, encompassing questionnaire-based variables (e.g., smoking, physical activity, fruit intake) and medical measurements (e.g., BMI, cholesterol levels), with a total of 21 features, to predict the presence of diabetes. The MNIST dataset is a comprehensive database of handwritten digits frequently used to benchmark image classification algorithms. Each image is a 28x28 pixel grayscale representation of a digit, ranging from 0 to 9. To increase task difficulty and highlight client differences in performance, our experiments utilise only 10,000 of the available 70,000 images, referred to as small-MNIST. Additionally, we transform these images into color, randomly assigning colors with equal probabilities: red, green, or blue. To explore the decision-making process through counterfactuals, we employed a ResNet-18 architecture to extract 1,000 features from each image.
@@
IID distributions across clients are achieved by randomly selecting samples from each dataset. In contrast, non-IID distributions are created using KMeans clustering [60] to form $K$ clusters of samples within each class—10 for the Synthetic dataset and 5 for others. We then assign the samples of the two nearest clusters from different classes to each client. To ensure each client’s distribution is equally represented, we randomly extract 15% of the samples from each client and combine them into the test set to assess the performance of our experiments. We create a clean validation set on the server by extracting 89 samples for Breast Cancer and 500 samples for other datasets. Additionally, we partition each client’s dataset locally, allocating 80% for training and 20% for validation.
@@ A.2 Model configuration
In our work, the model is composed of a predictor, any Deep Neural Network, which in our is a multi-layer perceptron that takes tabular data as input and makes a prediction. It is composed of five layers with hidden dimension equal to 512, 256, 256, 64, respectively. In addition, in our work, the model is also composed of a counterfactual generator, which can be implemented in different ways. The only requirement is that it can be trained concurrently with the main model, in an end-to-end fashion. To generalise our results, we tested the model with two different counterfactual generators. In Section 5, we used a counterfactual generator designed by us that explicitly optimises the generation of counterfactual with respect to specific labels using two VAEs, named CFGen and described in Appendix A.3. On the other hand, in Appendix B we also tested VCNet [34], which could generate counterfactual training a VAE, implementing it according to the original paper. In the context of the Federated Behavioral Space creation, PCA is employed as the dimensionality reduction technique $\psi_{n\to 2}$ within the EBP, centering the reduced space around a zero vector, which signifies the absence of errors. For CBP, tSNE is utilised as the dimensionality reduction method $\psi_{K\to 2}$, ensuring the preservation of distances between clients and their respective clusters. Finally, in terms of the implementation of Federated Behavioural Shields, at each round $t$, the score $s^{(k)}(t)$ for each client $k$ is computed using a moving average with a window length of $L$, as follow:
@@ A.3 CFGen Architecture
CFGen is a latent variable model that generates counterfactuals through variational inference. To this end, we have two random variables $z$ and $z^{\prime}$. These variables represent latent factors of variation whose probability distributions are easier to model and sample compared to those for $x$ and $x^{\prime}$. We also include dependencies from $y$ to the counterfactual latent distribution $z^{\prime}$ in order to explicitly model the dependency of $z^{\prime}$ on the class labels, resulting in the following overall probabilistic graphical model:
@@
Amortised inference. CFGen amortise inference needed for training by introducing two approximate Gaussian posteriors $q(z|\hat{x})$ and $q(z^{\prime}|z,\hat{x},y,y^{\prime})$ whose mean and variance are parametrised by a pair of neural networks $(\phi_{\mu},\phi_{\sigma})$ ($(\phi_{\mu^{\prime}},\phi_{\sigma^{\prime}})$, respectively).
Optimization problem. In order to obtain these counterfactuals, it is important to optimise their generation during training. CFGen is trained to maximise the log-likelihood of tuples $(\hat{x},y,y^{\prime})$, while observing $x$. Following a variational inference approach, we optimise the evidence lower bound of the log-likelihood, which results in the following objective function to maximise:
@@ A.4 Training configuration
Gradient Descent was employed as the optimisation algorithm, with a batch size equivalent to the dimension of the training dataset. Both the momentum and learning rate were set at 0.9 and 0.01, respectively. For centralised training scenarios, the model was trained over 1,000 epochs. The Flower library was utilised to implement FL [61]. In all federated experiments, except for those evaluating local epochs, we employed 2 local epochs. During the assessment of various defense mechanisms, the number of communication rounds was capped at 200 and the window length for the moving average of 30 rounds. Additionally, after evaluating the influence of test size on the server (see Appendix B), test sizes were set at 89 for Breast Cancer and 500 for the other datasets. For comparisons with centralised training, 1,000 rounds were used. For client personalization, the generator was trained across 25 local epochs. In each experiments, performance metrics were evaluated on the model that exhibited the lowest aggregated loss during training. The aggregated loss represents the weighted average of client losses, evaluated on each client’s local validation set, proportional to the respective number of samples.
@@ A.5 Code, licenses and hardware
For our experiments, we implement all baselines and methods in Python 3.9 and relied upon open-source libraries such as PyTorch 2.2 [62] (BSD license), Sklearn 1.4 [63] (BSD license), Flower 1.6 [61] (Apache License). In addition, we used Matplotlib [64] 3.8.2 (BSD license) and Seaborn [65] 0.13 (BSD license) to produce the plots shown in this paper. Data processing is performed using Pandas [66] 2.2 (BSD license). The three datasets we used are freely available on the web with licenses: Breast Cancer Wisconsin (CC BY-NC-SA 4.0 license), Diabetes Health Indicators (CC0 licence), and MNIST (GNU license).
We will publicly release the code with all the details used to reproduce all the experiments under an MIT license. All experiments were conducted on a workstation equipped with an NVIDIA RTX A6000 GPU, two AMD EPYC 7513 32-Core processors, and 512 GB of RAM.
A.6 Robust aggregation baselines

### v2

@@ 5.3 Leveraging FBPs information
FBPs offer detailed insights into client behaviours, enabling the Federated Behavioural Shields to outperform existing state-of-the-art defense mechanisms (Figure 4). Our comparative analysis demonstrates that Federated Behavioural Shields generally outperform traditional methods such as Krum, Median, and Trimmed-mean across various datasets including Breast, Diabetes, and small-MNIST. The only exception was in the scenario of the Inverted gradient on the small-MNIST dataset, where Trimmed-mean performed better. Overall, our method enhanced the performance up to 10 percentage points (pp) over Median—the most robust aggregation baseline—when the FL system is under Label-flipping attack and up to 16 pp when the system is not under attack on small-MNIST. The proposed approach surpasses even a FedAvg aggregation in absence of attackers on the Breast Cancer and on the small-MNIST datasets, under both normal and crafted-noise conditions. These results suggest that an aggregation strategy based on predictive performance or decision-making similarities is superior to methods that solely consider sample count. Unlike other baselines, the independence of our method from prior knowledge about attackers establishes it as a robust aggregation tool in both adversarial and non-adversarial settings. Contrary to previous studies [14, 28], we find that predictive performance alone does not reliably identify malicious clients in non-IID settings, as honest clients may also show consistently low performance, leading to potential misclassification in the aggregation process. Occasionally, this reliance on predictive performance can reduce our method’s accuracy compared to scenarios where only counterfactual information is used. For instance, under no-attack conditions, counterfactual information alone often offers more effective behavioural descriptors than combining it with predictive metrics. Further detailed analyses in Appendix B.12 indicate that using both descriptors generally yields better results than relying solely on counterfactual behaviour. Furthermore, visualizations in Appendix B.5 show how client scores consistently identify malicious clients during training, while Appendix B.11 demonstrates the robustness of our Federated Behavioural Shields against varying attack intensities.
@@ 6.1 Related works
Client behaviours in FL have been analysed using integrated visual tools [31, 46, 47] or indirectly through methods such as robust aggregation [10, 11, 12, 13, 14, 28, 30, 32, 40, 48, 49, 50, 51, 52] and clustering-based aggregation [53, 54, 55, 56, 57, 58, 59]. These studies typically focus on similarities in model or gradient parameters [10, 11, 12, 13, 31, 40, 46, 47, 48, 49, 50, 51, 52, 55, 56, 57, 58, 59], under the premise that distinct client data distributions manifest as unique model parameters. Nevertheless, these methods might overlook valuable information about how client models behave with the data, particularly in terms of predictive performance. To mitigate this issue, previous studies adopted evaluation methods that utilise a clean validation set on the server to delineate clients’ behaviour. Common descriptors used in such cases are straightforward metrics such as accuracy, error rates, and loss [14, 31, 46, 47, 53, 54]. Despite their utility, these metrics might not be able to reveal behavioural patterns, such as those involved in the decision-making processes of models, which may suggest subtle similarities or manipulations. Notably, Wang et al. [46] introduced a post-hoc explainable approach, Grad-CAM [60], to explain client model behaviours during training. However, Grad-CAM is limited to CNN models and provides primarily qualitative visual insights, which cannot be easily automated and thus still require human intervention. Similarly, SentiNet [32] uses Grad-CAM to detect potentially malicious regions in input images, yet client behaviour assessments continue to rely primarily on evaluating predictive performance on manipulated and unmanipulated images. To the best of our knowledge, this work represents the first systematic attempt to formalise the evolving dynamics of clients in FL, showing how behavioural client trajectories affect the predictive performance and decision-making processes of the global model.
@@ 6.2 Limitations and future works
The primary constraint of our framework lies in its assumption that the server possesses a minimal validation set for querying client models. While this assumption is common across various methods, it can be mitigated, as demonstrated in Wang et al. [46], by generating synthetic data points. To this end, our approach might already integrate a potential mechanism as differentiable counterfactual methods can be used to generate synthetic data [61]. Furthermore, as indicated by the promising results in Appendix B.6, utilizing validation-independent descriptors, such as counterfactuals, renders our defense method robust against biased or unfair validation sets. Given the significance of ensuring fairness across clients, developing additional validation-independent descriptors represents a promising direction for future research. Another consideration is the computational overhead introduced by counterfactual generators, which, although minimal compared to other baselines (see Appendix A.6), is higher than that of traditional FedAvg. However, this overhead can be mitigated by using a smaller network for the counterfactual generator, thereby reducing the number of neurons (e.g., 1.8% of predictor parameters) without compromising accuracy.
Future work could leverage the extensive information provided by FBPs to explore additional strategies for optimizing the learning process, such as the development of Clustered FL among clusters of clients and the fine-grained categorization of attack types. Incorporating additional behavioural planes may also enhance the specificity of the FBPs explanations. Lastly, since our method allows for the integration of privacy-enhancing techniques such as Local Differential Privacy [62] and Homomorphic Encryption [63], future studies could analyze their impact on the overall performance and computational efficiency of our system.
@@ 6.3 Conclusions
In this work, we proposed Federated Behavioural Planes, a method to explain the dynamics of FL systems and client behaviours. This innovative method allows to visualise, track, and analyse client behaviours based on specific characteristics. Our focus was twofold: examining predictive performance by analysing prediction errors and investigating the decision-making process through counterfactual generation. The results of our experiments showed that Federated Behavioural Planes enable to track client behaviours over time, cluster similar clients, and identify clients’ contributions to the global model with respect to a specific descriptor. Based on Federated Behavioural Planes’ information, we introduced a novel robust aggregation mechanism that improves existing state-of-the-art methods by not requiring prior knowledge of the attacker. This work lays the foundation to explain the evolution of client behaviours, with the potential to enhance reliability and control over FL systems.
Acknowledgments and Disclosure of Funding
This research was supported by the Swiss National Science Foundation, the European Union’s Horizon Europe research program, and the Slovenian Research and Innovation Agency under SmartCHANGE (No. 101080965), TRUST-ME (No. 205121L_214991), and XAI-PAC (No. Z00P2_216405) projects.
Author Contributions. Dario Fenoglio: Conceptualization, Methodology, Implementation, Experiments, Writing – Original Draft, Writing and supplementary experiments – Rebuttal and Final Paper. Gabriele Dominici: Conceptualization, Methodology, Implementation of Counterfactual Generators, Writing – Original Draft. Pietro Barbiero: Conceptualization, Review, Supervision. Alberto Tonda: Review. Martin Gjoreski: Review, Supervision, Funding Acquisition. Marc Langheinrich: Review, Supervision, Funding Acquisition.
@@ A.1 Datasets
In our experiments, we employ five distinct datasets:
  • Synthetic dataset (Tabular): It consists of two features randomly extracted from a range of -5 to +5. As shown in Figure 6.a, we partition this feature space into $K$ slices, where $K$ equals the number of clients. Each slice corresponds to a specific client. We assign a label of 1 to all data points where $x_{1}>\alpha x_{2}$, effectively creating a linear decision boundary that varies with $\alpha$. For each client, we generally draw 1000 samples. This controlled setup allows us to precisely manipulate and visually understand the data distribution across different clients, as illustrated in Figure 6a.
  • Breast Cancer Wisconsin (Tabular): It contains data from 569 patients with 30 continuous variables derived from digitised images of a fine needle aspirate of a breast mass. The variables describe characteristics of the cell nuclei present in the image (e.g., radius, area, perimeter), aimed at predicting breast cancer [36].
  • Diabetes Health Indicator (Tabular): It comprises data from 70,692 patients, encompassing questionnaire-based variables (e.g., smoking, physical activity, fruit intake) and medical measurements (e.g., BMI, cholesterol levels), with a total of 21 features, to predict the presence of diabetes [37].
  • MNIST (Image): It is a comprehensive database of handwritten digits frequently used to benchmark image classification algorithms [38]. Each image is a 28x28 pixel grayscale representation of a digit, ranging from 0 to 9. To increase task difficulty and highlight client differences in performance, our experiments utilise only 10,000 of the available 70,000 images, referred to as small-MNIST. Additionally, we transform these images into color, randomly assigning colors with equal probabilities: red, green, or blue. To explore the decision-making process through counterfactuals, we employed a ResNet-18 architecture to extract 1,000 features from each image.
  • CIFAR-10 (Image): The CIFAR-10 dataset comprises 60,000 color images categorised into 10 classes, widely used for evaluating object recognition algorithms [39]. Each image has a resolution of 32x32 pixels and is represented in RGB format. Similar to our small-MNIST setup, we reduce the sample size to 10,000 images. Features are extracted using a ResNet-18 architecture, resulting in 1,000-dimensional feature vectors for each image.
@@
IID distributions across clients are achieved by randomly selecting samples from each dataset. In contrast, non-IID distributions are created using KMeans clustering [64] to form $K$ clusters of samples within each class—10 for the Synthetic dataset and 5 for others. We then assign the samples of the two nearest clusters from different classes to each client. To ensure each client’s distribution is equally represented, we randomly extract 15% of the samples from each client and combine them into the test set to assess the performance of our experiments. We create a clean validation set on the server by extracting 89 samples for Breast Cancer and 250 samples for other datasets. Additionally, we partition each client’s dataset locally, allocating 80% for training and 20% for validation.
@@ A.2 Model configuration
In our work, the model is composed of a predictor, any Deep Neural Network, which in our is a multi-layer perceptron that takes tabular data as input and makes a prediction. It is composed of five layers with hidden dimension equal to 512, 256, 256, 64, respectively. In addition, in our work, the model is also composed of a counterfactual generator, which can be implemented in different ways. The only requirement is that it can be trained concurrently with the main model, in an end-to-end fashion. To generalise our results, we tested the model with two different counterfactual generators. In Section 5, we used a counterfactual generator designed by us, adapting the intuition from Dominici et al. [65], that explicitly optimises the generation of counterfactual with respect to specific labels using two VAEs, named CFGen and described in Appendix A.3. On the other hand, in Appendix B.2 we also tested VCNet [34], which could generate counterfactual training a VAE, implementing it according to the original paper. In the context of creating Federated Behavioural Planes, PCA is employed as the dimensionality reduction technique $\psi_{n\to 2}$ within the EBP, centering the reduced space around a zero vector, which signifies the absence of errors. For CBP, tSNE is utilised as the dimensionality reduction method $\psi_{K\to 2}$, ensuring the preservation of distances between clients and their respective clusters. Finally, in terms of the implementation of FBSs, at each round $t$, the score $s^{(k)}(t)$ for each client $k$ is computed using a moving average with a window length of $L$, as follow:
@@ A.3 CFGen architecture
CFGen is designed to adapt the approach proposed by  Dominici et al. [65] to tabular data, eliminating the need for predefined concepts. It is a latent variable model that generates counterfactuals through variational inference. To this end, we have two random variables $z$ and $z^{\prime}$. These variables represent latent factors of variation whose probability distributions are easier to model and sample compared to those for $x$ and $x^{\prime}$. We also include dependencies from $y$ to the counterfactual latent distribution $z^{\prime}$ in order to explicitly model the dependency of $z^{\prime}$ on the class labels, resulting in the following overall probabilistic graphical model:
@@
Amortised inference.
CFGen amortise inference needed for training by introducing two approximate Gaussian posteriors $q(z|\hat{x})$ and $q(z^{\prime}|z,\hat{x},y,y^{\prime})$ whose mean and variance are parametrised by a pair of neural networks $(\phi_{\mu},\phi_{\sigma})$ ($(\phi_{\mu^{\prime}},\phi_{\sigma^{\prime}})$, respectively).
Optimization problem.
In order to obtain these counterfactuals, it is important to optimise their generation during training. CFGen is trained to maximise the log-likelihood of tuples $(\hat{x},y,y^{\prime})$, while observing $x$. Following a variational inference approach, we optimise the evidence lower bound of the log-likelihood, which results in the following objective function to maximise:
@@ A.4 Training configuration
Gradient Descent was employed as the optimisation algorithm, with a batch size equivalent to the dimension of the training dataset. Both the momentum and learning rate were set at 0.9 and 0.01, respectively. For centralised training scenarios, the model was trained over 1,000 epochs. The Flower library was utilised to implement FL [66]. In all federated experiments, except for those evaluating local epochs, we employed 2 local epochs. During the assessment of various defense mechanisms, the number of communication rounds was capped at 200 and the window length for the moving average of 30 rounds. For comparisons with centralised training, 1,000 rounds were used. For client personalization, the generator was trained across 25 local epochs. In each experiments, performance metrics were evaluated on the model that exhibited the lowest aggregated loss during training. The aggregated loss represents the weighted average of client losses, evaluated on each client’s local validation set, proportional to the respective number of samples.
@@ A.5 Code, licenses and hardware
For our experiments, we implement all baselines and methods in Python 3.9 and relied upon open-source libraries such as PyTorch 2.2 [67] (BSD license), Sklearn 1.4 [68] (BSD license), Flower 1.6 [66] (Apache License). In addition, we used Matplotlib [69] 3.8.2 (BSD license) and Seaborn [70] 0.13 (BSD license) to produce the plots shown in this paper. Data processing is performed using Pandas [71] 2.2 (BSD license). The four datasets we used are freely available on the web with licenses: Breast Cancer Wisconsin (CC BY-NC-SA 4.0 license), Diabetes Health Indicators (CC0 license), MNIST (GNU license), and CIFAR-10. Our code, along with all the necessary details to reproduce the experiments, is publicly available on GitHub  [^3] https://github.com/dariofenoglio98/CF_FL under the MIT license. Additionally, we provide pseudo-code for both client-side (Algorithm 2) and server-side (Algorithm 1) implementations of our proposed approach, which includes creating behavioural planes on the server and applying our FBSs. All experiments were conducted on a workstation equipped with an NVIDIA RTX A6000 GPU, two AMD EPYC 7513 32-Core processors, and 512 GB of RAM.
Algorithm 1 The Federated Behavioural Shields Algorithm
1:Initial model $f(\theta(0))$, number of communication rounds $T$, number of selected clients per round $M$, clean validation set $(x^{(server)},y^{(server)})$, set of plane functions $S$
2:for $t=0,1,\ldots,T-1$ do
3:     Sample $M$ out of $N$ clients
4:     Send $\theta(t)$ to the selected $M$ clients
5:     wait
6:     for each selected client $k$ in $M$ in parallel do
7:         $x^{\prime(k)},y^{(k)}=f(\theta^{(k)}(t+1),x^{(server)})$
8:     end for
9:     for $plane\_fn$ in $S$ do $\triangleright$ Creation of planes
10:         $s_{j}^{k}(t)=plane\_fn(x^{\prime(k)},y^{(k)},x^{\prime(k)},y^{(server)})$triangleright Our case: Eq.5 and Eq.6-7
11:     end for
12:     Behavioural planes visualization (FBPs)
[…]
14:     $f(\theta(t+1))=\text{Model aggregation in Eq. 8 (part 2) with }s^{(k)}(t)$triangleright Eq. 2 for FBPs
15:end for
Algorithm 2 Local Training on Client $k$
1:Initial model architecture $f$, number of local epochs $E$, local dataset $(x^{(k)},y^{(k)})$
2:Receive current global model parameters $\theta(t)$ from the server
3:Initialise client model with global parameters: $f(\theta^{(k)}(t))\leftarrow f(\theta(t))$
4:Update local model $\theta^{(k)}(t+1)$ by training for $E$ epochs $\triangleright$ Both predictor and generator
5:Send updated model parameters $\theta^{(k)}(t+1)$ back to the server
A.6 Computational cost
This section outlines the computational costs associated with our proposed method, focusing on three primary components: local computation, communication overhead, and server-side computation.
Local computation.
Our methods (both FBPs and FBSs) integrate a counterfactual generator with the original predictor to analyse decision-making and provide insights into the client’s data distribution. For small neural networks, local computation is minimal compared to other costs in the FL framework, such as communication latency and synchronization. In these cases, the counterfactual generator can produce counterfactuals for the model input without impacting training efficiency. As the predictor size increases, as shown with MNIST and CIFAR-10, we can efficiently generate counterfactuals for intermediate layers using a relatively small number of neurons. For instance, in our experiments with small-MNIST and small-CIFAR-10, we used a ResNet-18, which has 12.42M parameters and requires 71.1M GFLOPs for inference with an input RGB image of size 28x28. As shown in Table 2, the generator with embedding size of 128 contributes only 5.1% of the operations compared to the predictor alone. Furthermore, by reducing the embedding size to 32, we maintained performance while reducing GFLOPs to just 2.7% of the ResNet-18.
Communication overhead.
Similar to the predictor, the counterfactual generator must be transmitted to the server for evaluation and aggregation, thereby increasing the number of model parameters sent. However, the additional communication overhead is marginal compared to the size of the predictor. As illustrated in Table 2, the counterfactual generator with an embedding size of 32, which maintains high task performance, consists of only 1.8% of the parameters of the ResNet-18. Consequently, with 32-bit precision, our counterfactual requires 0.92 megabytes (MB) out of approximately 49.68 MB for the predictor.
Server-side computation.
On the server, our methods involve evaluating client models’ performances on a clean validation set and calculating the pair-wise distances of the generated counterfactuals between clients. As shown in Figure 13, the validation set size can be relatively small, e.g., 250 samples, which can be processed with a single forward pass of the model. Consequently, the models’ evaluation on the server, which can also be parallelised, is negligible compared to the computational load of calculating pair-wise distances between client counterfactuals.
The primary computational bottleneck on the server is the pair-wise Wasserstein distance between client counterfactuals. To address this, we use the sliced Wasserstein distance implementation, which has a computational complexity of $O(m\log m)$, where $m$ is the number of supports, given by $m=\text{n\_samples}\times\text{reduction dimension ($\psi_{K\to 2}$)}$. In our implementation, $m=250\times 2=500$, and this operation is repeated for each unique pair of clients (i.e., binomial coefficient $\binom{n}{2}$), leading to a computational complexity of $O(n^{2}\cdot m\log m)$ where $n$ is the number of selected clients.
Compared to Krum, which has a complexity of $O(n^{2}d)$ with $d$ being the number of model parameters, our pair-wise operation is more efficient for neural networks with more than approximately 4480 parameters, which is likely the case in most practical applications.
Overall computational cost.
We evaluated the overall computational cost of our defense mechanism in comparison to both the Krum algorithm and the traditional FedAvg algorithm, considering different network sizes and varying numbers of clients. In all experiments, a validation set of 250 samples was used. For tests involving an increasing number of clients, our method incorporated a counterfactual generator utilizing 7.6% of the predictor’s parameters (in the worst-case scenario), as described in the original paper. Computational time was measured over 10 training rounds across 3 folds, and we reported the mean and standard error of the time per round. As shown in Figure 5, our method’s computational cost scales more efficiently with both the number of model parameters and the number of clients when compared to Krum. Specifically, our method adds only an additional minute per round for 200 clients compared to FedAvg, while Krum introduces over 15 minutes per round. Importantly, our approach maintains the robustness and security benefits, demonstrating its efficiency and scalability in large-scale FL scenarios.
Table 2: Comparison of embedding sizes of the counterfactual generator
Metrics Model Parameters GFLOP
Embedding size Accuracy Validity Pred.+CF CF Increase Pred.+CF CF Increase
128 (paper) $86.0\pm 0.3\%$ $100.0\pm 0.0\%$ $13.36M$ $0.94M$ $7.6\%$ $74.7M$ $3.6M$ $5.1\%$
64 $85.6\pm 0.4\%$ $100.0\pm 0.0\%$ $12.88M$ $0.47M$ $3.8\%$ $73.6M$ $2.5M$ $3.5\%$
32 $87.6\pm 0.5\%$ $100.0\pm 0.0\%$ $12.65M$ $0.23M$ $1.8\%$ $73.1M$ $1.9M$ $2.7\%$
((a))
((b))
Figure 5: Comparison of the computational time of our proposed method against the traditional FedAvg, and the robust aggregation Krum per round of training, across (a) different model parameters and (b) different numbers of clients.
A.7 Robust aggregation baselines